OpenStudy (anonymous):

A plane is 90 miles south and 110 miles east of London Heathrow Airport. What bearing should be taken to fly directly to the airport? Round your answer to one decimal place. Please HELP!!

OpenStudy (anonymous):

\[\text{Round}\left[270+\frac{180 \tan ^{-1}\left(\frac{90}{110}\right)}{\pi },0.1\right]=309.3 \text{ degrees}\]